OpenStudy (anonymous):

find range of the function 10-x^2

OpenStudy (anonymous):

f(x)=10-x^2 let f(x)=y so y=10-x^2

OpenStudy (anonymous):

then make x the subject x^2=10-y x= \[\sqrt{10-y}\] therefore f'(x)= \[\sqrt{10-y}\]

OpenStudy (anonymous):

and there you go :)

OpenStudy (anonymous):

understood?

OpenStudy (anonymous):

RANGE is fron 10 to infinity because x2 can never be negative.